What is the work for this..
y=-5x+17.
6x-y=16.
Answer is 3 for substitution.

Answer by richwmiller(17219)

Huh? Which answer? x or y?
6x-(-5x+17)=16
6x+5x-17=16
11x=33
x=3
y=-5(3)+17
y=2